# Time-varying Vector Field Compression with Preserved Critical Point Trajectories

### v2

@@ Abstract
I Introduction
Large-scale scientific simulations and observations produce massive amounts of data that overwhelm data storage and transmission systems. For instance, direct numerical simulation (DNS) of forced isotropic turbulence on the Frontier supercomputer [frontier] generates 0.5 PB of data in a single run [yeung2025small], and the ERA5 Atmospheric Reanalysis, which records climate data from 1979 to the present, has already accumulated to 5 PB [hersbach2020era5]. Such large data sizes not only make it difficult for producers to store the data but also hinder researchers and scientists from accessing it remotely.
Time-varying vector fields, such as flow velocity in DNS and atmospheric wind speed in ERA5, constitute a significant component of scientific data. These data describe how vectors change over both space and time, and are widely used to compute derived quantities and extract features of interest [gunther2017generic, friederici2015finite, fritschi2019visualizing, rimensberger2019visualization]. As such, enabling timely access to time-varying vector fields is a pressing need across many scientific and engineering disciplines. This often requires efficient data compression due to unprecedented data sizes.
[…]
@@ I Introduction
While error-controlled lossy compressors provide error guarantees over the raw data, most of them are agnostic to topological features. Topological features usually contain key information for novel scientific discoveries, and distortions in them may lead to wrong interpretations of the data and, thus, misleading discoveries. Critical point trajectories are important topological features of time-varying vector fields. They are time-parameterized curves for which the vector value equals zero at every point, and they are widely used for tracking cyclones in climatology [hewson2010objective], identifying eddies in oceanology [nencioli2010vector], and tracing vortices in computational fluid dynamics [BERSON200961, CoresofSwirling]. Given their role as primary analytical primitives, distortions in critical-point trajectories can lead to systematic errors in downstream analytics. For example, corrupted trajectories may result in incorrect detection of event initiation or termination, inaccurate estimation of feature lifetimes and movement patterns, and inconsistent query results when tracking features across time [TRICOCHE2002249, 1432684, fff]. Such errors can further propagate into misleading physical interpretations and flawed predictive models [neu2013imilast, woollings2018blocking, bai2020time, haller2005objective]. While prior works [liang2020toward, liang2022toward, xia2024preserving] enable the preservation of critical points during error-controlled lossy compression of single-snapshot data, they may still distort critical-point trajectories because the connections of critical points in space-time are overlooked.
@@
In this work, we propose the first highly effective lossy compression framework that exactly preserves critical-point trajectories for time-varying vector fields. Specifically, we extend the theory in [xia2024preserving] for preserving critical points in space to that of preserving critical-point trajectories in space-time, and then develop a lossy compression framework with exact critical-point trajectory preservation. To further improve compression efficiency, we propose a semi-Lagrangian predictor that better exploits spatiotemporal correlation in advection-dominated regions of time-varying vector fields, and combine it with the traditional Lorenzo predictor [ibarria2003out] for enhanced data decorrelation. To this end, we evaluate our framework on five real-world scientific datasets with six baselines. In summary, our contributions are as follows.
@@
[…]
The remaining sections of the paper are organized as follows. Section II discusses related works. Section III reviews the background on time-varying vector fields, critical-point trajectories, and the existing critical-point preserving compression algorithm. Section IV formulates the research problem. In Section V, we introduce the design of our compression framework with the preservation of critical-point trajectories. In Section VI, we present the semi-Lagrangian predictor and how we combine it with the Lorenzo predictor. In Section VII, we present and analyze the evaluation results. Finally, we conclude with a vision for future work in Section VIII.
@@ II-A Lossy Compression for Scientific Data
Traditional lossless compression, for example, LZ-family compressors such as gzip [gzip], Zstd [zstd], and LZ4 [lz4], as well as floating-point-oriented tools like FPZIP [fpzip], typically achieve a limited compression ratio on scientific data [lindstrom2017error]. This limitation has motivated the community to adopt error-bounded lossy compression to achieve a higher compression ratio while maintaining user-specified error metrics. Error-bounded lossy compressors usually follow the steps: decorrelation, quantization, and encoding. Depending on the decorrelation method, we can categorize these compressors into prediction-based compressors and transform-based compressors. The SZ compressor family [sz16, sz17, sz18, zhao2021optimizing] is a set of representative prediction-based compressors. They use predictors such as Lorenzo and regression for decorrelation and the decorrelated data are further quantized and encoded. Transform-based compressors leverage specific transforms for data decorrelation, and then perform quantization and encoding in the transformed domains. ZFP [lindstrom2014fixed] is a typical transform-based compressor that divides data into non-overlapped data blocks for independent compression. Other notable transform-based compressors include TTHRESH [ballester2019tthresh], SPERR [li2023lossy], and FAZ [liu2023faz]. MGARD [ainsworth2018multilevel, ainsworth2019multilevel, ainsworth2019qoi] is another lossy compressor that lies in the middle of prediction-based and transform-based compressors. It relies on multilinear interpolation with $L^{2}$ projection for data decorrelation, and applies level-wise quantization.
Beyond engineering practice, the community has also studied error propagation and numerical stability, including error analyses, statistical modeling of compression error distribution [lindstrom2017error]. Together, these results support replacing lossless compression with controlled-lossy approaches when application-level quantities of interest(QoIs) [pu2022qoi, liu2024qoi] remain within tolerance. In short, given the limited ratios of lossless methods on floating-point data, error-bounded lossy compression has become the mainstream solution for scientific data reduction and lays the groundwork for higher-level, structure-aware guarantees. This motivates topology-aware compressors that explicitly preserve structural invariants.
@@ II-B Topology-Aware Compression
Although error-bounded lossy compressors can enforce constraints such as $L^{\infty}$ or $L^{2}$ bounds, many scientific analyses care primarily about the topological structure of scalar or vector fields. For example, the number and locations of critical points, the connectivity of level sets like contours and isocontours, and the stability of contour/merge trees and Morse-Smale complexes. Purely amplitude-based error bounds can create or annihilate critical points or alter connectivity, undermining feature tracking and subsequent analysis.
For the scalar field data, a compressor building on persistent homology has been proposed [soler2018topologically]. This approach guarantees that high-persistence features are preserved exactly under a user-specified persistence threshold, while low-persistence noise is simplified and encoded. Methods such as TopoSZ [toposz] incorporate topological invariants directly into the prediction-quantization pipeline of SZ [sz-framework], preserving the contour tree of scalar fields while meeting pointwise error bounds. MSz and its derived variants [li2024msz, multitierMSZ] are capable of preserving the Morse-Smale complex structure while simultaneously leveraging parallelization to significantly reduce compression time. A complementary line of work [GeneralizedTopo] performs local or global topological simplification with $L^{\infty}$ control. Although this method is not compression algorithms per se, it can be integrated with existing compressor frameworks to preserve the topological structure of the data.
The need for topology-aware constraints naturally extends from scalar fields to vector and tensor fields. For example, the cpSZ framework [liang2022toward] is capable of preserving critical points during lossy compression, while cpSZ-SoS [xia2024preserving] further extends this framework into a parallelized compressor suitable for large-scale datasets. At the same time, it replaces the original numerical scheme with the Simulation of Simplicity (SoS) method for critical-point computation, thereby avoiding numerical instabilities that can occur in extreme cases. Building upon these foundations, subsequent work [xia2025tspsz] further extends such preservation to separatrix structures. More recently, the TFZ framework [TFZ] has introduced a topology-aware lossy compression strategy for 2D second-order tensor fields, illustrating that topology-aware compression is evolving from scalar fields toward higher-dimensional and more complex data types.
@@
In this section, we introduce the necessary background and preliminaries related to our work, including the notations, fundamental concepts, and techniques that form the basis of our proposed method.
@@ III-A Notation and Assumptions
We assume the target vector field is piecewise linear, which is a common practice in the community [nielson1991asymptotic, de2008computational, lorensen1987marching]. Following this assumption, the original domain is decomposed into different simplices, each of which can be represented by a linear function. Formally, let $D\subset\mathbb{R}^{2}$ be a polygonal domain endowed with a time-invariant simplicial mesh $\mathcal{M}=(\mathcal{V},\mathcal{E},\mathcal{F})$, where $\mathcal{V}$, $\mathcal{E}$, and $\mathcal{F}$ denote the sets of vertices, edges, and faces, respectively. The vertex positions are fixed in space across discrete times $\mathcal{T}=\{t_{0},\dots,t_{T}\}$. A time-varying 2D vector field is given by
@@ III-B Critical Point
A critical point is defined as a location at which the vector field vanishes. Under the piecewise-linear assumption over a 2D simplex mesh, a critical point can be computed by solving a linear system expressed in barycentric coordinates:
@@
where $(\mu_{0},\mu_{1},\mu_{2})$ are the barycentric coordinates, and $(u_{i},v_{i})$ denote the vector components at the vertices that constitute the cell. A critical point exists within the cell if and only if $0\leq\mu_{k}\leq 1$ holds for all $k\in\{0,1,2\}$. Figure  1 illustrates examples of cells in which a critical point exists and does not exist, respectively.
Figure 1: (a) Barycentric coordinates inside the cell(a critical point exists); (b) Barycentric coordinates outside the cell(a critical point does not exist)
@@ III-C Space–Time Extrusion and Tetrahedralization
Figure 2: (a) presents an example of the temporal evolution of the critical point and the corresponding events; (b) illustrates how lossy compression alters the critical point, leading to a change in its trajectory.
To address this issue, we adopt the method proposed by De Loera et al. [tetrahedron_split], which provides a unified representation of space and time by extruding the spatial domain along the temporal dimension, followed by a consistent space–time decomposition. At each time step $t_{k}$, we have a consistent triangular mesh $\Delta\in\mathcal{F}$. Then we extrude $\Delta$ over the temporal interval $t_{k},t_{k+1}]$ to form a triangular prism. This prism is then subdivided into three tetrahedra, denoted as $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$, using a fixed three-tetrahedron splitting scheme: $\tau_{1}=(\mathcal{V}_{t}^{a},\,\mathcal{V}_{t}^{b},\,\mathcal{V}_{t}^{c},\,\mathcal{V}_{t+1}^{c})$, $\tau_{2}=(\mathcal{V}_{t}^{a},\,\mathcal{V}_{t}^{b},\,\mathcal{V}_{t+1}^{b},\mathcal{V}_{t+1}^{c})$, and $\tau_{3}=(\mathcal{V}_{t}^{a},\,\mathcal{V}_{t+1}^{a},\,\mathcal{V}_{t+1}^{b},\mathcal{V}_{t+1}^{c})$. Within any tetrahedron $\tau\subset D\times[t_{k},t_{k+1}]$, the components $(u,v)$ are affine in $(x,y,t)$; hence the joint zero-set $\{\mathbf{V}=0\}$ is either empty or a straight line segment in $\tau$. As illustrated in Figure 3, this subdivision yields a refined and temporally coherent discretization of the data, ensuring a uniform topological structure across consecutive time steps. This consistency is crucial for the accurate detection and tracking of critical points in time-dependent vector fields.
@@
III-D Critical Point Trajectory
A critical point trajectory over $t_{k},t_{k^{\prime}}]$ is a polyline obtained by concatenating the line segments $\{\mathbf{V}=0\}$ across adjacent tetrahedra in $D\times[t_{k},t_{k^{\prime}}]$.
Under piecewise-linear assumption, the intersection $\{\mathbf{V}=0\}\cap(D\times\{t_{k}\})$ is governed by a face-level test: for a space–time triangular face $f=\{i,j,k\}$, if the origin lies in $\mathrm{conv}\{\mathbf{v}_{i},\mathbf{v}_{j},\mathbf{v}_{k}\}\subset\mathbb{R}^{2}$, then the zero-set crosses $f$ and yields a slice critical point.

### v3

@@ Abstract
Scientific simulations and observations are generating massive volumes of time-varying vector field data, posing significant challenges for long-term storage and data transmission. Lossy compression is widely regarded as a promising approach for reducing data volume, as lossless methods typically achieve only modest compression ratios and therefore provide limited reduction. However, directly applying existing lossy compression techniques to time-varying vector fields can introduce undesirable distortions in critical-point trajectories, which encode essential structural properties of the underlying field. In this work, we present an efficient lossy compression framework that exactly preserves all critical-point trajectories in time-varying vector fields. Our contributions are threefold. First, we extend the theory of critical-point preservation from the spatial domain to space-time and develop a corresponding compression framework to guarantee trajectory preservation. Second, we introduce a semi-Lagrangian predictor to more effectively exploit spatiotemporal correlations in advection-dominated regions, and integrate it with the classical Lorenzo predictor to further improve compression efficiency. Third, we evaluate the proposed approach against state-of-the-art lossy and lossless compressors on four real-world scientific datasets. Experimental results show that our method achieves compression ratios of up to $124.48\times$ while effectively preserving all critical-point trajectories. This compression ratio is up to $56.07\times$ higher than the best-performing lossless compressors. In contrast, none of the existing lossy compressors can preserve all critical-point trajectories at comparable compression ratios.
I Introduction
Large-scale scientific simulations and observations produce massive amounts of data that overwhelm data storage and transmission systems. For instance, direct numerical simulation (DNS) of forced isotropic turbulence on the Frontier supercomputer [frontier] generates 0.5 PB of data in a single run [yeung2025small], and the ERA5 Atmospheric Reanalysis, which records climate data from 1979 to the present, has already accumulated to 5 PB [hersbach2020era5]. Such large data sizes not only make it difficult for producers to store the data but also hinder researchers from accessing it remotely.
Time-varying vector fields, such as flow velocity in DNS and atmospheric wind in ERA5, constitute a significant component of scientific data. These data describe how vectors change over both space and time, and are widely used to compute derived quantities and extract features of interest [gunther2017generic, friederici2015finite, fritschi2019visualizing, rimensberger2019visualization]. As such, enabling timely access to time-varying vector fields is a pressing need across many scientific and engineering disciplines. This often requires efficient data compression due to unprecedented data sizes.
@@ I Introduction
While error-controlled lossy compressors provide error guarantees over the raw data, most of them are agnostic to topological features. Topological features usually contain key information for novel scientific discoveries, and distortions in them may lead to wrong interpretations of the data and, thus, misleading discoveries. Critical point trajectories are important topological features of time-varying vector fields. They are time-parameterized curves for which the vector value equals zero at every point, and they are widely used for tracking cyclones in climatology [hewson2010objective], identifying eddies in oceanology [nencioli2010vector], and tracing vortices in computational fluid dynamics [BERSON200961, CoresofSwirling]. Given their role as primary analytical primitives, distortions in critical-point trajectories can lead to systematic errors in downstream analytics. For example, corrupted trajectories may result in incorrect detection of event initiation or termination, inaccurate estimation of feature lifetimes and movement patterns, and inconsistent query results when tracking features across time [TRICOCHE2002249, 1432684, fff]. Such errors can further propagate into misleading physical interpretations and flawed predictive models [neu2013imilast, woollings2018blocking, bai2020time, haller2005objective]. While prior works [liang2020toward, liang2022toward, xia2024preserving] enable the preservation of critical points during error-controlled lossy compression of single-snapshot data, they may still distort critical-point trajectories, as the spatiotemporal connectivity of critical points is not explicitly considered.
@@
In this work, we present the first highly effective lossy compression framework that exactly preserves critical-point trajectories in time-varying vector fields. Specifically, we extend the theory in [xia2024preserving] from preserving critical points in the spatial domain to preserving critical-point trajectories in space-time, and develop a lossy compression framework that guarantees exact trajectory preservation. To further improve compression efficiency, we propose a semi-Lagrangian predictor that better exploits spatiotemporal correlation in advection-dominated regions, and combine it with the traditional Lorenzo predictor [ibarria2003out] for enhanced data decorrelation. To this end, we evaluate our framework on five real-world scientific datasets with six baselines.
In summary, our contributions are as follows.
@@
  • We evaluate the proposed method and compare it with three widely used lossless compressors and three leading lossy compressors using five scientific datasets. Experiments demonstrate that our method faithfully preserves all critical-point trajectories while delivering up to $56\times$ higher compression ratios than lossless compressors. In contrast, existing lossy compressors produce noticeable distortions in critical-point trajectories at similar or lower compression ratios.
The remainder of this paper is organized as follows. Section II reviews related work. Section III provides background on time-varying vector fields, critical-point trajectories, and existing critical-point-preserving compression algorithms. Section IV formulates the research problem. Section V presents the design of our compression framework for preserving critical-point trajectories. Section VI introduces the semi-Lagrangian predictor and describes its integration with the Lorenzo predictor. Section VII reports and analyzes the experimental results. Finally, Section VIII concludes the paper and outlines directions for future work.
@@ II-A Lossy Compression for Scientific Data
Lossless compression has been widely adopted in various domains such as database systems [hu2025icde, liu2025bittuner, lixi2024icde, tan2024dcc, tan2024ts]. Traditional lossless compression techniques, such as LZ-family compressors including gzip [gzip], Zstd [zstd], and LZ4 [lz4], floating-point-oriented tools like FPZIP [fpzip], typically achieve a limited compression ratio on scientific data[lindstrom2017error]. This limitation has motivated the community to adopt error-bounded lossy compression to achieve a higher compression ratio while maintaining user-specified error metrics. Error-bounded lossy compressors usually follow three steps: decorrelation, quantization, and encoding. Depending on the decorrelation method, we can categorize these compressors into prediction-based compressors and transform-based compressors. The SZ compressor and its derived variants [sz16, sz17, sz18, zhao2021optimizing, liu2025hpdc, wuxuan2025ipdps] represent a class of prediction-based compressors. They use predictors such as Lorenzo and regression for decorrelation and the decorrelated data are further quantized and encoded. Transform-based compressors leverage specific transforms for data decorrelation, and then perform quantization and encoding in the transformed domains. ZFP [lindstrom2014fixed] is a typical transform-based compressor that divides data into non-overlapped data blocks for independent compression. Other notable transform-based compressors include TTHRESH [ballester2019tthresh], SPERR [li2023lossy], and FAZ [liu2023faz]. MGARD [ainsworth2018multilevel, ainsworth2019multilevel, ainsworth2019qoi] is another lossy compressor that occupies a middle ground between prediction-based and transform-based approaches. It performs data decorrelation via multilevel multilinear interpolation with an $L^{2}$ projection, followed by level-wise quantization.
Beyond engineering practice, the community has also studied error propagation and numerical stability, including error analyses and statistical modeling of compression error distribution [lindstrom2017error]. Together, these results support replacing lossless compression with controlled-lossy approaches when application-level quantities of interest (QoIs) [pu2022qoi, liu2024qoi, wuxuan2025sc] remain within tolerance. In short, given the limited compression ratios of lossless methods on floating-point data, error-bounded lossy compression has become the mainstream solution for scientific data reduction and lays the groundwork for higher-level, structure-aware guarantees. This motivates topology-aware compressors that explicitly preserve structural invariants.
@@ II-B Topology-Aware Compression
Although error-bounded lossy compressors can enforce constraints such as $L^{\infty}$ or $L^{2}$ bounds, many scientific analyses are primarily concerned with the topological structures of scalar or vector fields [guoxi2025gale, guoxi2023atask]. Examples include the number and locations of critical points, the connectivity of level sets (i.e., contours), and the stability of contour trees, merge trees, and Morse-Smale complexes. Purely amplitude-based error bounds can create or annihilate critical points or alter connectivity, undermining feature tracking and subsequent analysis.
For the scalar field data, a compressor building on persistent homology has been proposed [soler2018topologically]. This approach guarantees that high-persistence features are preserved exactly under a user-specified persistence threshold, while low-persistence noise is simplified and encoded. Methods such as TopoSZ [toposz] incorporate topological invariants directly into the prediction-quantization pipeline of SZ [sz-framework], preserving the contour tree of scalar fields while meeting pointwise error bounds. An generalization of TopoSZ [nathan2025pvis] abstracts contour-tree preservation into a standalone layer that can be integrated with arbitrary lossy compressors. MSz and its derived variants [li2024msz, multitierMSZ] are capable of preserving the Morse-Smale complex structure while simultaneously leveraging parallelization to significantly reduce compression time. A complementary line of work [GeneralizedTopo] performs local or global topological simplification with $L^{\infty}$ control. Although this method is not compression algorithms per se, it can be integrated with existing compressor frameworks to preserve the topological structure of the data.
The need for topology-aware constraints naturally extends from scalar fields to vector and tensor fields. For example, the cpSZ framework [liang2022toward] is capable of preserving critical points during lossy compression, while cpSZ-SoS [xia2024preserving] further extends this framework into a parallelized compressor suitable for large-scale datasets. At the same time, it replaces the original numerical scheme with the Simulation of Simplicity (SoS) method for critical-point computation, thereby avoiding numerical instabilities that can occur in extreme cases. Building upon these foundations, subsequent work [xia2025tspsz] further extends such preservation to separatrix structures. More recently, TFZ [TFZ] introduces a topology-aware lossy compression strategy for 2D second-order tensor fields, illustrating that topology-aware compression is evolving from scalar fields toward higher-dimensional and more complex data types.
@@
In this section, we present the necessary background and preliminaries for our work, including the notation, fundamental concepts, and techniques that underpin the proposed method.
@@ III-A Notation and Assumptions
We assume the target vector field is piecewise linear, which is a common practice in the community [nielson1991asymptotic, de2008computational, lorensen1987marching]. Following this assumption, the original domain is decomposed into simplices, each of which can be represented by a linear function. Formally, let $D\subset\mathbb{R}^{2}$ be a polygonal domain endowed with a time-invariant simplicial mesh $\mathcal{M}=(\mathcal{V},\mathcal{E},\mathcal{F})$, where $\mathcal{V}$, $\mathcal{E}$, and $\mathcal{F}$ denote the sets of vertices, edges, and faces, respectively. The vertex positions are fixed in space across discrete times $\mathcal{T}=\{t_{0},\dots,t_{T}\}$. A time-varying 2D vector field is given by
@@ III-B Critical Point
A critical point is defined as a location where the vector field vanishes. Under the piecewise-linear assumption over a 2D simplicial mesh, a critical point can be computed by solving a linear system expressed in barycentric coordinates:
@@
where $(\mu_{0},\mu_{1},\mu_{2})$ are the barycentric coordinates, and $(u_{i},v_{i})$ denote the vector components at the vertices that constitute the cell. A critical point exists within the cell if and only if $0\leq\mu_{k}\leq 1$ holds for all $k\in\{0,1,2\}$. Figure 1 illustrates examples of cells with and without a critical point, respectively.
Figure 1: (a) Barycentric coordinates inside the cell, where a critical point exists; (b) barycentric coordinates outside the cell, where no critical point exists.
@@ III-C Space–Time Extrusion and Tetrahedralization
Figure 2: (a) illustrates the temporal evolution of a critical point and the associated events; (b) demonstrates how lossy compression alters the critical point, resulting in a deviation of its trajectory.
To address this issue, we adopt the method proposed by De Loera et al. [tetrahedron_split], which provides a unified representation of space and time by extruding the spatial domain along the temporal dimension, followed by a consistent space-time decomposition. At each time step $t_{k}$, we have a consistent triangular mesh $\Delta\in\mathcal{F}$. Then we extrude $\Delta$ over the temporal interval $t_{k},t_{k+1}]$ to form a triangular prism. This prism is then subdivided into three tetrahedra, denoted as $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$, using a fixed three-tetrahedron splitting scheme: $\tau_{1}=(\mathcal{V}_{t}^{a},\,\mathcal{V}_{t}^{b},\,\mathcal{V}_{t}^{c},\,\mathcal{V}_{t+1}^{c})$, $\tau_{2}=(\mathcal{V}_{t}^{a},\,\mathcal{V}_{t}^{b},\,\mathcal{V}_{t+1}^{b},\mathcal{V}_{t+1}^{c})$, and $\tau_{3}=(\mathcal{V}_{t}^{a},\,\mathcal{V}_{t+1}^{a},\,\mathcal{V}_{t+1}^{b},\mathcal{V}_{t+1}^{c})$.
Within any tetrahedron $\tau\subset D\times[t_{k},t_{k+1}]$, the components $(u,v)$ are affine in $(x,y,t)$; hence the joint zero-set $\{\mathbf{V}=0\}$ is either empty or a straight line segment in $\tau$. As illustrated in Figure 3, this subdivision yields a refined and temporally coherent discretization of the data, ensuring a uniform topological structure across consecutive time steps. This consistency is crucial for the accurate detection and tracking of critical points in time-dependent vector fields.
@@
III-D Critical-Point Trajectory
A critical-point trajectory over $t_{k},t_{k^{\prime}}]$ is a polyline obtained by concatenating the line segments $\{\mathbf{V}=0\}$ across adjacent tetrahedra in $D\times[t_{k},t_{k^{\prime}}]$.
Under piecewise-linear assumption, the intersection $\{\mathbf{V}=0\}\cap(D\times\{t_{k}\})$ is governed by a face-level test: for a space-time triangular face $f=\{i,j,k\}$, if the origin lies within the convex hull $\mathrm{conv}\{\mathbf{v}_{i},\mathbf{v}_{j},\mathbf{v}_{k}\}\subset\mathbb{R}^{2}$, then the zero set intersects $f$, yielding a slice critical point.